OpenStudy (anonymous):

what is the sum of the first 10 terms of the sequence defined by a^n=2^n-3

OpenStudy (anonymous):

\[a_{n}=2n-3\]

OpenStudy (anonymous):

a1=2*1-3=-1,a2=2*2-3=1 n=10 c.d d=1-(-1)=1+1=2 n=10 l=2*10-3=17 \[Sn=\frac{ n }{2}\left( a1+l \right)\] plug the values and get S10

OpenStudy (anonymous):

thanks!